Q1.) An aerial camera has a 9½-inch square film format and a 6-inch focal length lens. What must the flying height (in meter) be to obtain a scale of 1/2000? (Hint: 1 inch=2.54 cm)

Q2.)If the distance of satellite on elliptical orbit at the point of A from two focal points is 3AU and 1AU on 21st of June, what would be the sum of distance of satellite from point B to two focal points on 21st of December in astronomical unit:

Q3.) For a spacecraft travelling from Earth to Mars what would happen if the rocket motor is turned off following the initial speed required to enter into a hohmann transfer orbit?
Answer #1

Ans 1 : The flying height is 304.8 meters.

Given :

Aerial camera has 9½-inch square film format

Focal length : 6 inch

Air photo scale : 1/2000

To find : Flying height ( H )

Solution :

Air photo scale at a particular point is Sp = f / (H-h),

where f is the focal length of the camera,

H is the flying height of the aircraft above mean sea level,

and h is the elevation of the surface at a given point.


Sp = f /( H-h );
H-h = f / Sp

h = 0 ( No elevation of surface )

Convert Focal length f in meters = 6 inches x 2.54cm/inch = 15.24 cm = 0.1524 m

H = f/ Sp

H = 0.1524 m/1/2000

= 304.8 meters

Hence, the flying height is 304.8 meters.